# On the origin of circular rolls in rotor-stator flow

### v1

@@ 7 Conclusions and outlooks
The present numerical investigation has focused on closed axisymmetric rotor-stator flow as a prototype flow for the study of receptivity. Motivated by the riddle about the origin of the dynamical circular rolls observed in several experiments (Savas, 1987; Gauthier, 1998; Schouveiler et al., 2001), both linear and nonlinear receptivity theories were invoked to investigate how the flow responds to an additive forcing of external origin. Several types of forcing were considered, each with its pros and cons. Forcing applied to the bulk of the fluid lends itself well to linear optimal response theory (Schmid & Brandt, 2014). If a wide frequency spectrum is forced, pseudo-resonance linked with non-normal effects (Trefethen & Embree, 2005) will amplify more the frequencies contained in the large hump in the gain curve in fig. 9. Such frequencies invariably lead to circular rolls inside the Bödewadt layer developing along the stator. Forcing applied through motion of the boundaries is by design not optimal, however it is more realistic in the case of closed flows, and can easily be extended to a nonlinear framework. Increasing the amplitude of the forcing at high enough $Re$ unambiguously leads to a different, more complex nonlinear state which coexists with the simpler forced base flow. Although both states are characterised by circular rolls on the stator, we argue that the rolls reported in the experiments of Schouveiler et al. (2001) are a linear response to external forcing and not a self-sustained state like those found by Gesla et al. (2023). This is consistent with the observation that these rolls disappear rapidly should the forcing be suddenly turned off (Lopez et al., 2009; Do et al., 2010). Moreover a particularly striking match with the experiment of Schouveiler et al. (2001) was found when the forcing features only harmonics of the rotor’s main angular frequency, with complex features such as the pairing and merging of vortices well reproduced numerically. The circular rolls observed experimentally can hence be described as noise-sustained, with the subtlety that they are not the response to incoherent noise but rather to a forcing with well-organised frequency content typical of rotating machinery.
@@
As far as the larger picture is concerned, although linear receptivity is the starting point for transition theory based on an input-output picture, any theory aiming at predicting the onset of bistable dynamics needs to be nonlinear. Extending the resolvent formalism to incorporate nonlinear interactions is a possibility (Rigas et al., 2021). As is the case in the present study, the optimality idea from classical linear resolvent theory gets lost in the nonlinear framework in favour of other nonlinear selection criteria. A generalisation of the concept of state space, laminar-turbulent boundary, edge states and other finite-amplitude states to forced problems is awaited for, and seems possible given the encouraging results in figure 16. A clearer connection between the input-output formalism (Sipp & Marquet, 2013), former theories with nonlinear selection criteria in spatially developing flows of global modes (Pier & Huerre, 2001) and the dynamical systems framework would be welcome.

### v2

@@ 7 Conclusions and outlooks
The present numerical investigation has focused on closed axisymmetric rotor-stator flow as a prototype flow for the study of receptivity. Motivated by the riddle about the origin of the dynamical circular rolls observed in several experiments (Savas, 1987; Gauthier et al., 1999; Schouveiler et al., 2001), both linear and nonlinear receptivity theories were invoked to investigate how the flow responds to an additive forcing of external origin. Several types of forcing were considered, each with its pros and cons. Forcing applied to the bulk of the fluid lends itself well to linear optimal response theory (Schmid & Brandt, 2014). If a wide frequency spectrum is forced, pseudo-resonance linked with non-normal effects (Trefethen & Embree, 2005) will amplify more the frequencies contained in the large hump in the gain curve in fig. 9. Such frequencies invariably lead to circular rolls inside the Bödewadt layer developing along the stator. Forcing applied through motion of the boundaries is by design not optimal, however it is more realistic in the case of closed flows, and can easily be extended to a nonlinear framework. Increasing the amplitude of the forcing at high enough $Re$ unambiguously leads to a different, more complex nonlinear state which coexists with the simpler forced base flow. Although both states are characterised by circular rolls on the stator, we argue that the rolls reported in the experiments of Schouveiler et al. (2001) are a linear response to external forcing and not a self-sustained state like those found by Gesla et al. (2024). This is consistent with the observation that these rolls disappear rapidly should the forcing be suddenly turned off (Lopez et al., 2009; Do et al., 2010). Moreover a particularly striking match with the experiment of Schouveiler et al. (2001) was found when the forcing features only harmonics of the rotor’s main angular frequency, with complex features such as the pairing and merging of vortices well reproduced numerically. The circular rolls observed experimentally can hence be described as noise-sustained, with the subtlety that they are not the response to incoherent noise but rather to a forcing with well-organised frequency content typical of rotating machinery.
The present study is fully axisymmetric, which implies that only axisymmetric modes are involved in the generation of the circular rolls. This excludes other scenarios involving non-axisymmetric modes, for instance triadic instabilities featuring the interaction of a given non-axisymmetric mode with its complex conjugate. A possible extension of our work would allow for non-axisymmetry already at the linear level. The resulting resolvent matrix would then depend on an additional azimuthal wavenumber $m$ and the optimally forced structures would have to be identified in a 2D parameter space $(\omega,m)$. When the time integration of the forced system is concerned a 3D time integration code would be necessary to study the response of the flow to the boundary forcing. While it is theoretically possible that 3D structures are more amplified that the axisymmetric ones, the experimental (Gauthier et al., 1999) and numerical (Do et al., 2010) evidence suggest that the response of the forcing is in the form of circular rolls, supporting the axisymmetric assumption.
@@
As far as the larger picture is concerned, although linear receptivity is the starting point for transition theory based on an input-output picture, any theory aiming at predicting the onset of bistable dynamics needs to be nonlinear. Extending the resolvent formalism to incorporate nonlinear interactions is a possibility (Rigas et al., 2021). A clearer connection between theories of nonlinear global modes in spatially developing flows (Pier & Huerre, 2001) and the dynamical systems framework would also be welcome.
Declaration of Interests : The authors report no conflict of interest.
Acknowledgement : The authors wish to thank anonymous referees for suggestions, in particular the superexponential fit of the lifetimes.